# GenCO: Generating Diverse Solutions to Design Problems with Combinatorial Nature

### v1

@@ Abstract
Generating diverse objects (e.g., images) using generative models (such as GAN or VAE) has achieved impressive results in the recent years, to help solve many design problems that are traditionally done by humans. Going beyond image generation, we aim to find solutions to more general design problems, in which both the diversity of the design and conformity of constraints are important. Such a setting has applications in computer graphics, animation, industrial design, material science, etc, in which we may want the output of the generator to follow discrete/combinatorial constraints and penalize any deviation, which is non-trivial with existing generative models and optimization solvers. To address this, we propose GenCO, a novel framework that conducts end-to-end training of deep generative models integrated with embedded combinatorial solvers, aiming to uncover high-quality solutions aligned with nonlinear objectives. While structurally akin to conventional generative models, GenCO diverges in its role - it focuses on generating instances of combinatorial optimization problems rather than final objects (e.g., images). This shift allows finer control over the generated outputs, enabling assessments of their feasibility and introducing an additional combinatorial loss component. We demonstrate the effectiveness of our approach on a variety of generative tasks characterized by combinatorial intricacies, including game level generation and map creation for path planning, consistently demonstrating its capability to yield diverse, high-quality solutions that reliably adhere to user-specified combinatorial properties.
@@ 1 Introduction
Generating diverse and realistic objects with combinatorial properties is an important task with many applications. For example, in video game level design, we may want to generate a variety of levels that are both realistic and valid/playable (Zhang et al., 2020). Here, “valid” may refer to certain discrete characteristics of the level that must be satisfied (e.g., a minimum number of enemies, a path between the level entrance and exit, etc.). In automatic device design, we want to generate a variety of devices that meet foundry manufacturing constraints and optimize physics-related objectives (Schubert et al., 2022). In the design of new molecules, we want to generate a variety of chemically valid molecules with specific properties (Pereira et al., 2021). In all of these examples, the goal is to generate a variety of combinatorial solutions that are both feasible and high quality with respect to a given nonlinear objective. Another situation to consider is when the generated solutions do not have to strictly follow combinatorial constraints, but we want to discourage certain things as a penalty. For example, if we are creating various images of a map, we might want to ensure that the path planning in those maps is efficient (see section 4.2).
The difficulty in these settings arises from the fact that the generated objects need to satisfy specific discrete properties incorporated either as combinatorial constraints or penalties and be of high quality as evaluated by a nonlinear objective. The combinatorial constraints make it challenging to apply gradient-based methods to train standard generative models, such as GAN (Goodfellow et al., 2014), or VAE (Kingma & Welling, 2013), which may generate examples that violate the user-specified constraints. Additionally, the nonlinear objective makes it difficult to use traditional general-purpose combinatorial optimization tools since mixed-integer nonlinear programming (MINLP) is often slow in practice compared to the more heavily researched mixed integer linear programming (MILP), often requiring problem-specific solvers or leveraging specialized structure. Furthermore, traditional optimization solvers are not geared towards generating diverse solutions and do not have the flexibility of deep learning models to quickly adapt to slightly modified settings. Recent work has attempted to address these issues by first training a generative model and then fixing/postprocessing the output after the fact (Zhang et al., 2020). However, these approaches are limited in that they only use the combinatorial solver to fix the output of the generative model and do not use the combinatorial solver in training. As a result, the generative model is not trained with the combinatorial solver in mind (i.e., end-to-end). Thus, the generator may fail to capture the distribution of high-quality feasible solutions as many of the generated examples may be “fixed” to a single solution. Additionally, some work constrains generative model outputs by penalizing the generator based on constraint violation (Chao et al., 2021). While this approach may work in some settings where the constraints must be satisfied, it is unclear how to extend this approach to more complex constraints, such as logical or general combinatorial constraints that can be readily expressed in MILP.
As a result, we propose GenCO that generates diverse solutions to design problems with combinatorial nature. GenCO combines the flexibility of deep generative models with the combinatorial efficiency of optimization solvers. Our approach mainly deals with two cases: (1) when the resulting solution needs to satisfy certain hard combinatorial constraints (hard constraints), and (2) the imposition of a desired penalty if the solutions violate the constraints (soft constraints). To achieve these two goals, we introduce a general concept known as the “combinatorial loss”, which serves as an additional “restriction” on the generated solutions. In the context of combinatorial constraints, the combinatorial loss takes the form of a projection expressed as a mixed-integer linear program (MILP). In the second scenario, this general concept manifests as an extra penalty term within the objective function. This, coupled with the diversity mechanisms inherent in deep generative models, fine-tunes the generated solutions. This process both encourages diversity and ensures adherence to the specified combinatorial properties.
Our main contributions involve introducing a framework, GenCO, that serves two key purposes: 1) generating a range of solutions to nonlinear problems, ensuring they adhere to combinatorial properties; 2) integrating a combinatorial solver into the deep generative learning process. We showcase the effectiveness of our approach through a series of experiments on various generative combinatorial optimization problems. This includes tasks like game level design (section 4.1) and generating maps for path planning (section 4.2).
@@ 2 Related Work
The interaction between generative models and combinatorial optimization has seen increased research interest as practitioners seek to integrate or draw inspiration from both paradigms. A prominent line of research has approached the problem of generating objects with combinatorial optimization in mind.
@@
Further work has endeavored to modify deep generative models such as generative adversarial networks (Goodfellow et al., 2014) to penalize constraint violation. General purpose methods are proposed for generating objects respecting constraint graphs (Para et al., 2021) and blackbox constraints (Di Liello et al., 2020). Specialized approaches consider graph constraints where the generated object should be a graph that meets certain constraints that are specified by a blackbox function such as in the design of photonic crystals (Christensen et al., 2020), crystal structure prediction (Kim et al., 2020), graph-constrained house generation (Tang et al., 2023), and general house plan generation (Nauata et al., 2021). In these settings, generating infeasible objects is penalized, and the generative model often has an inductive bias in the network architecture to generate feasible objects. These methods can use historical training datasets and handle unstructured objects; however, the output is often infeasible.
@@
In a related direction, previous work proposed conditioning variational autoencoders (VAE) (Kingma & Welling, 2013) with combinatorial programs (Misino et al., 2022). Here, the latent information is extracted by the encoder and then fed through a logic program, made differentiable via DeepProblog (Manhaeve et al., 2018), such as performing handwritten operations on digits. The result is then fed through the decoder to generate the original image based on logical relationships. At test time, the goal is to generate objects that hopefully satisfy the logical relationship because the model was conditioned to do so. Here, the logical program helps to condition the generative model to generate objects based on patterns resulting from logical relationships and penalize cases where the logic is incorrect. Here, the generated objects are not guaranteed to satisfy the combinatorial constraints, but rather, the generative model has a logical structural prior. Previous work has made various optimization problems differentiable, such as quadratic programs (Amos & Kolter, 2017), probabilistic logic programs (Manhaeve et al., 2018), linear programs (Wilder et al., 2019a; Mandi & Guns, 2020; Elmachtoub & Grigas, 2017; Liu & Grigas, 2021), Stackelberg games (Perrault et al., 2020), normal form games (Ling et al., 2018), kmeans clustering (Wilder et al., 2019b), maximum likelihood computation (Niepert et al., 2021), graph matching (Rolínek et al., 2020), knapsack (Demirovic et al., 2019b; a), maxsat (Wang et al., 2019), mixed integer linear programs (Mandi et al., 2020; Paulus et al., 2021; Ferber et al., 2020), blackbox combinatorial solvers (Pogančić et al., 2020; Mandi et al., 2022; Berthet et al., 2020), nonlinear programs (Donti et al., 2017), continuous constraint satisfaction (Donti et al., 2020), cone programs (Agrawal et al., 2019). General-purpose methods are presented for minimizing downstream regret by learning surrogate loss functions (Shah et al., 2022; 2023; Zharmagambetov et al., 2023). Additionally, previous work has employed learnable linear solvers to solve nonlinear combinatorial problems (Ferber et al., 2023). Finally, a survey investigates the intersection between machine learning and optimization (Kotary et al., 2021). These approaches focus on identifying a single solution rather than generating diverse solutions. Additionally, many of these approaches are amenable for use by GenCO if a specific optimization problem better suits the generative problem.
@@
Recently, generative models have been proposed to improve combinatorial optimization. Zhang et al. (2022) use generative flow networks (gflownet) for robust scheduling problems. Sun & Yang (2023) use graph diffusion to solve combinatorial problems on graphs. Ozair et al. (2021) use vector quantized variational autoencoders to compress the latent space for solving planning problems in reinforcement learning. Zhao & You (2020) use generative adversarial networks to generate settings for sample average approximation in robust chance-constrained optimization. Additionally, in (Lopez-Piqueres et al., 2023), the authors generate continuous objects with tensor networks that satisfy linear constraints for optimization problems. These approaches target solving a fully specified optimization problem rather than generating objects with combinatorial constraints in mind.
@@ 3 GenCO: Method Description
Figure 1: An overview of GenCO. The top part of the diagram corresponds to the ordinary “GAN-style” generative model (chosen as an example). By integrating the “combinatorial loss” into both the loss function and the object generation process (bottom part), we effectively encode the combinatorial nature of the problem.
3.1 Mathematical Formulation
The key distinction in our framework centers around the incorporation of a combinatorial loss, a departure from conventional deep generative models like Generative Adversarial Networks (GANs) illustrated in Figure 1. Rather than directly generating the ultimate object of interest, be it images or any other complex entity, our generator $G{}(\epsilon{};\theta{})$ parameterized by $\theta{}$ takes random noise $\epsilon{}$ and transforms it into a problem representation $c$ that encapsulates the underlying problem’s essential features. This representation serves as a pivotal intermediary step. Subsequently, we feed $c$ to both the combinatorial loss and the generator loss (e.g. derived via a discriminator), ensuring that the solutions produced by the generator exhibit the desired characteristics: diversity, realism, and adherence to combinatorial properties.
To train such a generator, we define the following bi-objective function, effectively balancing the generative loss $\mathcal{L}_{\text{gen}}{}$ with the combinatorial loss $\mathcal{C}{}$:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(h(c))]$ (1)
s.t. $c=G(\epsilon;\theta{}),$
where $\mathcal{L}_{\text{gen}}{}$ encapsulates the generative loss, responsible for promoting diversity and realism. Its form can vary substantially depending on the specific generative model employed. For example, a Wasserstein GAN might utilize the adversarial discrepancy between outputs, while a Variational Autoencoder (VAE) would make use of the Evidence Lower Bound (ELBO).
Conversely, the combinatorial loss $\mathcal{C}{}(\cdot)$, penalized by the weighting factor $\gamma$, is coupled with the downstream optimization solver $h(\cdot)$. This solver plays a crucial role in enforcing or encouraging combinatorial properties. For instance, one might seek to generate objects that must adhere to physical or logical constraints, all while minimizing a specified (linear) cost. In such a scenario, it becomes imperative to formulate an instance of a Mixed-Integer Linear Program (MILP) or another type of optimization problem with a well-defined feasible region. This problem instance then integrates seamlessly into the training pipeline, serving as the backbone of $\mathcal{C}{}$.
Below, we consider two potential and practical settings for $\mathcal{C}{}$, each tailored to address specific real-world applications.
3.2 Constrained Generator
In this setting, we consider a problem domain where the goal is to train a generator that outputs feasible solutions, which natively handles diversity. To generate objects that are guaranteed to be feasible, we set $\gamma\to+\infty$ in equation 1 for any $c$ that does not satisfy constraints (defined by feasible region $\Omega$), 0 otherwise. To handle this directly with such a high value of $\gamma$ is computationally inefficient. Instead, we employ a ”projection layer” denoted as $h(\cdot)$ that maps $c$ to its nearest feasible point in the feasible set $\Omega$. Specifically, we define $h(c)=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$  [^1] Although a dot product is a natural choice of proximity for cosine distance, one could employ other metrics, e.g. $|c-x|$, where $c$ is an output of the generator $G(\epsilon;\theta)$. In this context, a generated problem description $c$ (as in Fig.1) can be viewed as an unconstrained generated object that is subsequently mapped via $h(\cdot)$ to $x$ to adhere to the defined constraints. With this approach, we can reformulate equation 1 into the following unconstrained optimization problem:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(h(G(\epsilon;\theta{})))].$ (2)
Note that the feasibility within region $\Omega$ is guaranteed to be satisfied due to the form of $h(\cdot)$. Directly minimizing this objective is possible thanks to the recent advancements in differentiable solvers literature (Agrawal et al., 2019; Pogančić et al., 2020; Sahoo et al., 2022).
Algorithm 1 provides a high-level description of our method for this setting. It starts by initializing the generator parameters. In each iteration, a problem description is randomly selected using the generator. An optimization problem is then solved to find the solution that maximizes the similarity between $c$ within a defined feasible set ($x$ is the corrected/feasible version of $c$). Finally, the generator parameters are updated through backpropagation. We also provide an application of this generic algorithm to the concrete generative model examples of GANs and VQVAEs in Appendices B-C.
@@ 3.2 Constrained Generator
  Initialize generator parameters $\theta$;
@@
     Sample a noise $\epsilon$;
     Sample a problem description $c\sim G(\epsilon;\theta)$;
     Call a solver $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x=h(c)$;
     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(x^{*})$;
     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
@@
Algorithm 1 Generator Training in Constrained Setting
3.3 Penalized Generator
The formulation from the previous section 3.2 naturally extends to penalty form, i.e., we encourage generated problem representation $c$ to respect certain combinatorial properties. In this scenario, we can directly leverage equation 1 as there is no need in projection layer:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta{}))+\gamma\mathcal{C}{}(h(G(\epsilon;\theta{})))].$ (3)
Here, we seamlessly integrate the generator into the objective function. While it is true that we do not impose any explicit constraints on the output of the generator, and one could use the output of $G$ directly, we contend that introducing a combinatorial penalty or loss as a penalty remains a crucial framework with significant practical utility. This approach serves to further guide the generation process and injects domain-specific knowledge that is easy to interpret, ensuring that the generated solutions exhibit desired combinatorial properties, a facet that holds immense practical relevance.
For instance, we may aim to generate images tailored for path planning in strategy games. These maps feature various types of terrains, each incurring distinct costs when placed on the map. For efficient path planning, our goal is to find a “low cost” shortest path between two endpoints. Therefore, in this scenario, $\mathcal{C}$ incorporates a domain specific (fixed) mapping (e.g. a neural net) that transforms an image into a graph with corresponding weights assigned to each edge, followed by solving the shortest path problem (e.g. using Dijkstra’s algorithm).
However, in this particular case, there is no need to introduce a “projection” operator as in the previous case since there are no explicit constraints that must be strictly satisfied. We can directly employ gradient-based optimization, allowing for differentiation through the combinatorial problem. This is achieved approximately by employing methods outlined in (Pogančić et al., 2020; Sahoo et al., 2022). Pseudocode 2 describes the application of GenCO for penalty setting. Furthermore, Appendix A.1 provides an application of this generic algorithm to a concrete example of GANs.
  Output: Trained generator
  Initialize generator parameters $\theta$;
  (optional) Define a fixed mapping $g(\cdot)$ for $c$
  while not converged do
     Sample a noise $\epsilon$;
     Sample a problem description $c\sim G(\epsilon;\theta)$;
     (optional) apply a mapping $g(\cdot)$ to transform $c$;
     Call a solver $x^{*}=h(g(c))$ ;
     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(x^{*})$;
     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
  end while
Algorithm 2 Generator Training in Penalty Setting
It is crucial to emphasize that the given penalized setting deviates from the conventional decision-focused learning (DFL) or smart predict-then-optimize paradigms (Elmachtoub & Grigas, 2017; Donti et al., 2017). In our approach, we do not possess access to the ground truth costs $c$, and the ultimate loss function encompasses both the generator’s loss and the output of a solver.

### v2

# GenCO: Generating Diverse Designs with Combinatorial Constraints

@@ Abstract
Deep generative models like GAN and VAE have shown impressive results in generating unconstrained objects like images. However, many design settings arising in industrial design, material science, computer graphics and more require that the generated objects satisfy hard combinatorial constraints or meet objectives in addition to modeling a data distribution. To address this, we propose GenCO, a generative framework that guarantees constraint satisfaction throughout training by leveraging differentiable combinatorial solvers to enforce feasibility. GenCO imposes the generative loss on provably feasible solutions rather than intermediate soft solutions, meaning that the deep generative network can focus on ensuring the generated objects match the data distribution without having to also capture feasibility. This shift enables practitioners to enforce hard constraints on the generated outputs during end-to-end training, enabling assessments of their feasibility and introducing additional combinatorial loss components to deep generative training. We demonstrate the effectiveness of our approach on a variety of generative combinatorial tasks, including game level generation, map creation for path planning, and photonic device design, consistently demonstrating its capability to yield diverse, high-quality solutions that verifiably adhere to user-specified combinatorial properties.
Machine Learning, ICML
[…]
usc,cornell,FAIR
@@ 1 Introduction
Generating diverse and realistic objects with combinatorial properties is an important task with many applications. For AI-guided design in particular, the goal is to generate a diverse set of combinatorial solutions that are both feasible and high quality with respect to a given nonlinear objective. Diversity is often useful when dealing with uncertain metrics, diverse preferences, or for the sake of further manual scrutiny. For example, in photonic device design, we want to generate a variety of devices that meet foundry manufacturing constraints and optimize physics-related objectives (Schubert et al., 2022). In the design of new molecules, we want to generate a variety of chemically valid molecules with specific properties (Pereira et al., 2021). In other scenarios, we would like to create a diverse set of designs so that their solutions (not necessarily diverse given the design) meet certain objectives. For example, in video game level design, we may want to generate levels that are both realistic and valid/playable (Zhang et al., 2020). Here, “valid” may refer to discrete characteristics of the level that must be satisfied (e.g., a minimum number of enemies, a guaranteed path between the level entrance and exit, etc.).
Approaches Solution Feasibility Data Nonlinear
generation guarantees distribution objective
GAN/VAE (Goodfellow et al., 2014; Kingma & Welling, 2013) Fast
Integer Programming (Deaton & Grandhi, 2014) Slow
Lagrangian Multiplier (Di Liello et al., 2020) Slow
SurCo (Ferber et al., 2023) Slow
Fixing/Rejection (Zhang et al., 2020; Gomes et al., 2006) Slow
GenCO (proposed) Fast
Table 1: A comparison of relevant directions tackling generative modeling or combinatorial optimization for design problems.
Mathematically, we seek several solutions $\mathcal{X}:=\{{\bm{x}}_{j}\}$ to the following optimization problem:
$\min_{\mathcal{X}}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})\quad\quad\mathrm{s.t.}\ {\bm{x}}_{j}\in\Omega$ (1)
where $\mathcal{L}$ is a group loss defined on a population of solutions $\mathcal{X}$ (e.g., minimizing the Wasserstein distance to a dataset), $\mathcal{D}$ is an individual loss measuring the quality of each solution ${\bm{x}}_{j}$, $\gamma$ is the weighing coefficient and $\Omega$ is the feasible set. Eqn. 1 can be solved by direct optimization or learning approaches that leverage solution datasets.
Several approaches can attempt to solve this problem. The Lagrangian (Frangioni, 2005) can reformulate the problem as $\min_{\mathcal{X}}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})+\lambda\sum_{j}\mathrm{dist}_{\Omega}({\bm{x}}_{j})$, where $\mathrm{dist}_{\Omega}({\bm{x}}_{j})$ measures the distance between ${\bm{x}}_{j}$ and the feasible set $\Omega$ ($\mathrm{dist}_{\Omega}({\bm{x}}_{j})=0$ if ${\bm{x}}_{j}\in\Omega$). However the final solutions ${\bm{x}}_{j}$ may not be feasible unless we use $\lambda\rightarrow+\infty$ which may lead to numerical instability. To reduce the optimization cost, we want to learn a generative model to directly generate ${\bm{x}}_{j}$, and avoid on-the-fly optimization. However, standard methods like GAN (Goodfellow et al., 2014) or VAE (Kingma & Welling, 2013) fail to generate feasible and high quality solutions. One approach (Chao et al., 2021) penalizes the generator based on constraint violation but is nontrivial to extend to more complex constraints, such as logical or general combinatorial constraints that can be readily expressed in MILP. Alternative approaches include postprocessing solutions (Zhang et al., 2020), which may lead to repetitive solutions since the generative network is not trained with the postprocessing, or rejection sampling (Gomes et al., 2006) that can be computationally inefficient due to high rejection rates. Fundamentally, the training procedure does not involve verifiably feasible solutions. Another approach is to leverage integer linear programming (ILP) to find feasible solutions to problems with linear constraints on integer variables. However, such solvers are not geared towards generating diverse solutions and cannot leverage deep learning models to capture the data distribution or quickly adapt to slightly modified settings.
In this work, we propose GenCO. Instead of directly optimizing ${\bm{x}}$, we learn a latent design space $\mathcal{C}$ so that (1) each element ${\bm{c}}\in\mathcal{C}$ yields a guaranteed feasible solution ${\bm{x}}\in\Omega$, often through a combinatorial solver ${\bm{g}}$: ${\bm{x}}={\bm{g}}({\bm{c}})$, and (2) a generative model $G_{\theta}$ can map noise to the latent space $\mathcal{C}$. The entire pipeline is trained in an end-to-end manner via backpropagation through the solver ${\bm{g}}$. Intuitively, this latent space $\mathcal{C}$ represents the hidden parameters that lead to the solution. For example, ${\bm{c}}$ represents the layout of a 10x10 maze, while ${\bm{x}}$ represents the optimal path of the maze.
GenCO addresses many challenges in previous work. GenCO generates solutions efficiently due to its learned generative model $G_{\theta}$. Additionally, GenCO guarantees feasibility via re-parameterization in the latent design space, which is different from conventional deep generative models like Generative Adversarial Networks (GANs) that directly generate the ultimate object of interest ${\bm{x}}$, be it images or any other complex entity. GenCO encourages diverse solutions, thanks to the group loss term $\mathcal{L}$ imposed on the solution population and end-to-end training of the entire pipeline. Finally, GenCO does not require linear objective functions and can handle nonlinear objectives. A comparison is summarized in  Table 1.
Our main contributions involve introducing a framework, GenCO, that can: 1) train generative models to generate solutions that are guaranteed to satisfy combinatorial properties throughout training; 2) optimize nonlinear objectives; 3) flexibly handle a variety of generative models and optimizers. We showcase the effectiveness of our approach empirically on various generative combinatorial optimization problems: creating game level ( subsection 4.2), generating maps for path planning ( subsection 4.3), and designing inverse photonic devices ( subsection 4.4).
@@ 2 Related Work
The interaction between generative models and combinatorial optimization has drawn attention as practitioners seek to generate objects with combinatorial optimization in mind.
@@
Further work has modified deep generative models such as GAN (Goodfellow et al., 2014) to penalize constraint violation. These methods can generate objects respecting constraint graphs (Para et al., 2021), and blackbox constraints (Di Liello et al., 2020). Specialized approaches consider blackbox-constrained graphs such as in the design of photonic crystals (Christensen et al., 2020), crystal structure prediction (Kim et al., 2020), and house generation (Tang et al., 2023; Nauata et al., 2021). Here, feasibility is encouraged through penalties and inductive bias but is not guaranteed.
Figure 1: GenCO diagram. A neural generator projects noise to a latent space which then gets used by a solver to generate provably feasible solutions. The solutions are then penalized with generative losses like Wasserstein distance (WGAN) or reconstruction (VQVAE) which can be backpropagated through the full pipeline. Additionally, GenCO can optimize individual objectives.
@@
Previous work proposed conditioning VAE (Kingma & Welling, 2013) with combinatorial programs (Misino et al., 2022). Here, the latent information is extracted by the encoder and then fed through DeepProblog, a differentiable logic program (Manhaeve et al., 2018). The result is then fed through the decoder to generate the original image based on logical relationships. At test time, the model generates objects that hopefully satisfy the logical relationship through model conditioning; however, there is no guarantee of constraint satisfaction. Previous work has made various optimization problems differentiable, such as quadratic programs (Amos & Kolter, 2017), probabilistic logic programs (Manhaeve et al., 2018), linear programs (Wilder et al., 2019a; Mandi & Guns, 2020; Elmachtoub & Grigas, 2017; Liu & Grigas, 2021), Stackelberg games (Perrault et al., 2020), normal form games (Ling et al., 2018), kmeans clustering (Wilder et al., 2019b), maximum likelihood computation (Niepert et al., 2021), graph matching (Rolínek et al., 2020), knapsack (Demirovic et al., 2019b, a), maxsat (Wang et al., 2019), mixed integer linear programs (Mandi et al., 2020; Paulus et al., 2021; Ferber et al., 2020), blackbox combinatorial solvers (Pogančić et al., 2020; Mandi et al., 2022; Berthet et al., 2020), nonlinear programs (Donti et al., 2017), continuous constraint satisfaction (Donti et al., 2020), cone programs (Agrawal et al., 2019). General-purpose methods are presented for minimizing downstream regret by learning surrogate loss functions (Shah et al., 2022, 2023; Zharmagambetov et al., 2023). Additionally, previous work has employed learnable linear solvers to solve nonlinear combinatorial problems (Ferber et al., 2023). Finally, there is a survey on machine learning and optimization (Kotary et al., 2021). These approaches focus on identifying a single solution rather than generating diverse solutions. Additionally, many of these approaches can be integrated into GenCO if a specific optimization problem better suits the generative problem.
@@
Recently, generative models have been proposed to improve combinatorial optimization. (Zhang et al., 2022) use generative flow networks (gflownet) for robust scheduling problems. (Sun & Yang, 2023) use graph diffusion to solve combinatorial problems on graphs. (Ozair et al., 2021) use vector quantized variational autoencoders to compress the latent space for solving planning problems in reinforcement learning. (Zhao & You, 2020) use generative adversarial networks to generate settings for sample average approximation in robust chance-constrained optimization. Additionally, in (Lopez-Piqueres et al., 2023), the authors generate continuous objects with tensor networks that satisfy linear constraints for optimization problems. These approaches solve fully specified optimization problems rather than generating objects with combinatorial constraints.
@@ 3 GenCO: Method Description
Experiment Sol ${\bm{x}}$ Feas $\Omega$ Group Loss $\mathcal{L}$ Ind. Loss $\mathcal{D}$ Latent ${\bm{c}}$ Solver ${\bm{g}}$
Game Design Game Level Playability WGAN (comb) Soft sol. Gurobi (ILP)
Path Planning Min Path Path WGAN Min path (comb) RGB map Gurobi (LP)
Photonic Device 0/1 Grid Manufacturing VQVAE (comb) Maxwell’s sim. (comb) Soft sol. Domain spec.
Table 2: Summary of experimental settings highlighting the broad applicability of GenCO to settings with various optimizers, generative models, and loss functions. Losses operating on combinatorial objects rather than continuous latent space are highlighted with (comb).
In the GenCO framework ( Figure 1), instead of direct optimization of the solution ${\bm{x}}$, we learn a distribution over a latent design space $\mathcal{C}$ so that for any ${\bm{c}}\in\mathcal{C}$, we get a solution ${\bm{x}}={\bm{g}}({\bm{c}})\in\mathcal{X}$ that is guaranteed to be feasible. Here ${\bm{g}}(\cdot)$ is a combinatorial solver that uses its input ${\bm{c}}$ to specify the combinatorial optimization problems that it solves. For example, ${\bm{g}}({\bm{c}}):=\operatorname*{arg\,min}_{{\bm{x}}\in\Omega}{\bm{c}}^{\top}{\bm{x}}$ outputs the solution to a mixed integer linear program (MILP) with coefficients ${\bm{c}}$. ${\bm{g}}(\cdot)$ can also output a shortest path (via Dijkstra’s algorithm), given a weighted graph represented by ${\bm{c}}$, etc.
During the training, we learn a generative model $G_{\theta}$ so that ${\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$ leads to solutions that minimize the loss:
$\min_{\mathcal{C}:=\{{\bm{c}}_{j}\}}\mathcal{L}(\{{\bm{g}}({\bm{c}}_{j})\})+\gamma\sum_{j}\mathcal{D}({\bm{g}}({\bm{c}}_{j}))\quad\mathrm{s.t.}\ {\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$ (2)
Here $\epsilon_{j}$ is random noise driving generation. This can be written as the following (with $\mathcal{X}:=\{{\bm{x}}_{j}\}$ representing a group of generated solutions):
$\min_{\theta}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})\quad\mathrm{s.t.}\ {\bm{x}}_{j}={\bm{g}}(G_{\theta}(\epsilon_{j}))$ (3)
The generative model parameters $\theta$ can be learned via backpropagation through the combinatorial solver ${\bm{g}}(\cdot)$.
Choices of group loss $\mathcal{L}$. The group loss $\mathcal{L}$ encourages the generated examples to capture the data distribution and can be instantiated using the Wasserstein distance as in Wasserstein GAN (Arjovsky et al., 2017), or Evidence Lower Bound (ELBO) as in VQVAE (Van Den Oord et al., 2017). In the WGAN case, the adversarial discriminator is trained only on examples satisfying the combinatorial constraints (${\bm{x}}$) rather than continuous solutions. Similarly, in VQVAE, the known feasible solutions are passed through the autoencoder followed by the solver ${\bm{g}}$ to ensure that the autoencoder loss is only computed on feasible solutions ${\bm{x}}$ rather than the continuous latent space ${\bm{c}}$.
Choices of individual loss $\mathcal{D}$. We design the individual loss $\mathcal{D}$ to control the quality of individual solutions (e.g., the manufacture cost of a generated molecule ${\bm{x}}_{j}$ should be minimized). In many settings, this individual loss is nontrivial to encode in combinatorial solvers like Gurobi as in (Ferber et al., 2023).
Loss on solution space $\Omega$ vs latent space $\mathcal{C}$ Note that to optimize GenCO’s objective ( Equation 3), we would need to run the combinatorial solver at every iteration, which involves calling MILP solvers and can be slow. On the other hand, if the group loss $\mathcal{L}$ or individual loss $\mathcal{D}$ can be evaluated in the latent space $\mathcal{C}$, then this extra cost is not needed during training. This occurs in the map generation experiments ( subsection 4.3), in which an observation ${\bm{c}}_{i}$ is provided (as an RGB image) and the optimal solution ${\bm{x}}_{i}$ can be computed accordingly (e.g., via an edge cost estimator and shortest path solver):
$\min_{\theta}$ $\mathcal{L}(\mathcal{C})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})$ (4)
$\ \ \mathrm{s.t.}$ $\ {\bm{x}}_{j}={\bm{g}}({\bm{c}}_{j})\quad\text{ and }\quad{\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$
This formulation is related to SurCo-zero (Ferber et al., 2023), in which a latent representation ${\bm{c}}$ is optimized so that its inferred solution ${\bm{x}}={\bm{g}}({\bm{c}})$ by the solver ${\bm{g}}$ minimizes a nonlinear loss function $\mathcal{D}$. Mathematically, the training objective of SurCo is $\min_{\bm{c}}\mathcal{D}({\bm{g}}({\bm{c}}))$. When there are additional problem descriptions ${\bm{y}}_{i}$ (i.e., SurCo-prior), the latent design vector ${\bm{c}}_{i}={\bm{c}}_{\theta}({\bm{y}}_{i})$ becomes a function of ${\bm{y}}_{i}$ and thus the SurCo objective becomes $\min_{\theta}\sum_{i}\mathcal{D}({\bm{g}}({\bm{c}}_{\theta}({\bm{y}}_{i}));{\bm{y}}_{i})$. The main difference here is that GenCO considers the group loss $\mathcal{L}$ (e.g., matching the data distribution) on newly generated solutions, while SurCo focuses on the quality of solutions to known problem descriptions during training.
In addition to  Figure 1, we summarize our proposed method in Algorithm 1. In each iteration, a set of solutions is generated using generator $G_{\theta}$ and a solver ${\bm{g}}$ (i.e., forward pass). We then evaluate the loss function using equation 3 (or equation 4). The parameters of $G_{\theta}$ are updated via backpropagation. Here, the differentiation through the combinatorial solver is achieved approximately by employing methods outlined in (Pogančić et al., 2020; Sahoo et al., 2022).
Assumptions
GenCO has several assumption on the individual loss $\mathcal{D}$, group loss $\mathcal{L}$, solver $g$, and latent space $c$. We assume the individual loss $\mathcal{D}$ is a function that maps a fixed solution to a loss value and provides gradient information useful for gradient descent. Note that if gradient information is not available, a neural network estimator may be learned to approximate the individual loss function. The group loss $\mathcal{L}$ imposes a loss on the distribution of instances generated. With deep generative models, this group loss is generally imposed by using learning-based proxies that act on individual examples. The assumption overall is that $\mathcal{L}$ can be optimized via stochastic gradient descent on the generated examples. We assume the solver $g$ is a function that maps from generated latent information $c$ to a feasible solution $x\in\Omega$. Additionally, we assume that previous work has made $g$ differentiable as a function of $c$ with various techniques proposed such as those described in related work. Finally, we assume the latent space $c$ simply represents values that can be mapped to via a neural network, and that the solver is well defined on that input.
@@
  Initialize generator parameters $\theta$
@@
     for $j=1\ldots K$ do
        Sample a noise $\epsilon_{j}$
        Sample a latent description ${\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$
        Call a solver ${\bm{x}}_{j}={\bm{g}}({\bm{c}}_{j})$
     end for
     Compute loss using  Equation 3 with $\mathcal{X}:=\{{\bm{x}}_{j}\}$ and $\mathcal{C}:=\{{\bm{c}}_{j}\}$.
     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$
@@
Algorithm 1 GenCO